# White Paper and Roadmap for Quantum Gravity Phenomenology in the Multi-Messenger Era

### v2

@@ Abstract
The unification of quantum mechanics and general relativity has long been elusive. Only recently have empirical predictions of various possible theories of quantum gravity been put to test. The dawn of multi-messenger high-energy astrophysics has been tremendously beneficial, as it allows us to study particles with much higher energies and travelling much longer distances than possible in terrestrial experiments, but more progress is needed on several fronts.
@@ 1 Introduction
To bridge this gap between fundamental theories of QG and observations, a bottom-up approach emerged: QG phenomenology [9]. It is a framework that employs phenomenological models inspired by the results from the different fundamental approaches and by physical reasoning about QG, to identify effects which can be searched for in experiments. How and which of these effects emerge from the theory, depends on the theoretical model under consideration [10, 11, 9, 12, 13], which helps devise the experimental setups required for searching for specific signatures, and the types of effects being sought. At the same time, evidence or bounds on any given effect are serving as guides for the construction of fundamental theories compatible with these phenomenological results.
@@
In this white paper the QG phenomenology community summarizes and highlights paths to advance and improve the search for QG effects.
The overall purpose of this white paper is to outline the future phenomenological and experimental research of QG with astrophysical data. We address some of the most promising observational windows to scales where we expect QG effects to set in, as well as the necessary conditions for realistically boosting these searches with present and future facilities. Even more importantly, we aspire to stimulate developments of novel methods for searches for signatures of QG. Given that at this point there are no facilities designed for this specific purpose, and experimental tests are usually performed as side projects, we particularly hope to motivate a development of dedicated experiments, and that this white paper can be used as a road map in such endeavours.
@@ 2.1.2 Some specific effects
A competing theoretical mechanism that can cause a cosmic vacuum refractive index is associated with certain exotic models of dark matter [47, 48]. They can be dinstingusihed from spacetime foam models through the specific dependence of the refractive index on the photon frequency and the emergence or non-emergence of birefringence [49].
@@
Birefringence (different propagation speeds for each light polarization) from QG spacetime with a polymer-like structure at microscales in loop quantum gravity [57]. Such effect is predicted by mass-dimension five operators in the effective field theory extension of the QED sector of the Standard Model [58]. Birefringence effects are further discussed in section 3.2.1.
@@ 2.2 Time delays in the propagation of cosmic messengers
Time delays emerge already from the classical interaction between gravity and particles and fields that is described by the general relativistic geometry of spacetime. In this section, we briefly recall this classical achromatic time delay effects, that are already predicted by GR, before examining additional energy-dependent (generally four-momentum dependent) time delays, due to microscopic quantum fluctuations or foaminess of spacetime (see section 2.1), which are widely discussed in the literature of QG and its phenomenology. Then, we highlight the main challenges and difficulties in the search for such QG-induced time delays.
@@ 2.2.1 Classical time delays
Even in the absence of QG effects, the gravitational interaction causes time delays in the propagation of photons of any frequency and other cosmic messengers through spacetime, compared to their propagation in the absence of gravity. The most famous instance of this fact, which has been experimentally observed and actually constitutes one of the Solar-System tests of Einstein’s GR, is the Shapiro delay [82]: the observation that the travel time of a radar signal between two points in the vicinity of a gravitating object is increased compared to the travel time of the signal in the absence of the object. It is important to stress that the Shapiro delay is “achromatic”, in the sense that it does not depend on the energy of the photon. In general, we gather large amounts of information about the gravitational interaction based on the observation of cosmic messengers propagating through spacetime and eventually being detect by us. Thus, in the pertinent data analyses of telescopes observing cosmic messengers, it is of utmost importance to take into account the classical – and possibly the yet unknown additional QG-induced – time delay effects, in order to interpret the observations correctly. Famous examples where the classical time delays are already implemented in the data analysis are pulsar timing observations [83, 84] as well as cosmological redshift and distance observations [85]. Pictorially, the time delay effects can be understood in the sense that gravity acts as a certain effective refractive medium for light propagation, thus introducing an effective refractive index [86]. On the basis of Maxwell’s electrodynamics and Einstein’s GR, the predictions for the classical time delays are independent of the energy (frequency) of the photons, as a consequence of the Einstein equivalence principle [87], which states that locally, that is, within an appropriately small region in the neighbourhood of each spacetime point, physics is described by special relativity (SR). In what follows we refer to this feature as local Lorentz symmetry. Due to the universal coupling with gravity, such GR-induced time delays do not only affect the propagation of light, but also all other standard model particles and fields including neutrinos [88] and, due to its self interaction, gravitational waves [89, 90].
The above examples from classical gravity clearly demonstrate that (achromatic) time delays are an important prediction of GR, which, on the one hand, allows us to test the theory, while on the other, gives us important information about the nature of the (classical) gravitational interaction, which is associated with the geometry of spacetime.
@@ 2.2.2 Possible quantum gravity-induced time delays
In addition to the general relativistic, achromatic time delays discussed so far, time delays resulting from purely quantum effects of gravity might emerge. In contrast to the classical effects, they might depend on the energy of the probe, for the the following reason. Highly energetic particles and fields will interact with gravity on smaller length scales than low energetic ones. It is expected that the smaller the length scale which the particles and fields probe, i.e. the higher their energy, the more the particles and fields are affected by quantum vacuum fluctuations of gravity or spacetime foam, see also the discussion in section 2.1. Effectively, this scenario can be described as particles and fields propagating through a (refractive) medium generated by vacuum fluctuations of gravity, capturing aspects of a quantum spacetime or spacetime foam. The mathematical description of this pictorial idea has been further developed within various formal and phenomenological frameworks of QG.
@@ 2.2.3 Theoretical challenges
A particularly important challenge is the distinction between violation and deformation of local Lorentz invariance. In other words, either the LIV implies the existence of a preferred reference frame, or the Lorentz group is replaced by the notion of spacetime and observer symmetry (DSR). Both, the absence or the detection of any kind of time delay can be used as a guiding principle to identify the scale and the type of deviations from local Lorentz symmetry. Whether this happens effectively at an intermediate length/energy scale or at the fundamental QG scale is unclear and depends on the theoretical approach under consideration.
@@
In summary, it is known that gravity causes certain types of time delays. The search for additional, yet undetected energy-dependent time delays caused by potential QG effects can be an important step in finding evidences of a theory of QG. Thus, it is crucial to devise future dedicated experiments to address such a possibility both in an astrophysical context and in that of the more controllable Earth-based experiments. These two regimes complement each other, in the sense that they probe QG time delays on different scales, which have different systematic uncertainties.
@@ 2.3 Modifications of the kinematics of interactions of cosmic messengers
Threshold anomalies can be seen as kinematic modifications of the SR description of interactions, and indeed, it is usually the only effect that is modelled to estimate the phenomenological consequences of LIV or DSR on the interactions of cosmic messengers [107, 94]. However, the decay rates or cross sections of a process in LIV or DSR scenarios may differ from the standard ones even at energy ranges where the process is allowed in both, SR and its modifications, which may be important for effects such as the possible modification of the transparency of the universe to high-energy gamma rays. The LIV scenario has been explored in the past, but the availability of results for cross sections or decay rates is limited [108, 109]. For the DSR scenario, changes in the cross sections or decay rates may only play a role in the case of a QG energy scale much lower than the Planck one, comparable to the energies of cosmic messengers. This is a mostly unexplored case, for which it will be necessary to develop new approaches and techniques [106].
@@ 2.3.1 Hadronic processes (affecting cosmic rays)
[…]
@@ 3.2.1 Propagation effects in vacuum
The temporal distribution of emitted particles is usually described with a light curve template defined by several parameters. The magnitude of uncertainties of these parameters varies from case to case, and is a matter of compromise between several factors. Faster and more pronounced flux variability puts stronger constraints on the emission time: the sensitivity to the QG energy scale is inversely proportional to the variability time scale for linear () QG effects and to its square root for quadratic () ones [94, Table 1]. This, however, requires using shorter time bins, which results in relatively smaller number of events per bin, and, consequently, larger uncertainties. For this reason, brighter, more variable emissions are selected to study time delays. Improvement of instrumental sensitivity will directly reflect on the ability to model the emission light curves, and on the sensitivity to the QG energy scale.
@@
[…]
@@ 4.3 Cosmic-ray experiments
Presently our most powerful cosmic-ray telescope is the Pierre Auger Observatory, and its results have provided the basis for several QG-phenomenology analyses [239, and references therein]. However, in general converting UHECR data into constraints on QG parameters faces several challenges due to our lack of knowledge about the sources and propagation of UHECR, described in section 3.1, but one can be optimistic about progress in these directions in the coming years [178].
@@ 6 Final remarks
The scale of is therefore one of the most pressing theoretical unknowns, and experimental efforts in QG phenomenology should aim for the identification of this scale. As a generic goal, this corresponds to trying to push the high-energy limit in the present and near-future experiments. This may be crucial for the observation of sharp cutoffs in the spectrum of photons or neutrinos predicted in superluminal LIV models, as well as to observe an increase of the transparency of the universe for high-energy photons or cosmic rays in the case of DSR or subluminal LIV models. Another relevant effect that may be observed with better exposition at high energies is an enhancement in the cosmogenic production of neutrinos and photons by the interactions of cosmic rays with the CMB and EBL.
@@
A.F. Ali (Essex County College, NJ, USA); M. Arzano (University of Naples and INFN); J-L. Atteia (IRAP, Université de Toulouse, CNRS, CNES, UPS, Toulouse, France); C. Barceló (Instituto de Astrofísica de Andalucía, IAA CSIC); I. Ben-Dayan (Ariel University, Israel); D. Benisty (University of Cambridge, and Frankfurt Institute for Advanced Studies, fias); N. Bilic (Rudjer Boskovic Institute, Zagreb, Croatia); O. Birnholtz (Bar Ilan University, Israel); D. Blas (”Institut de Física d’Altes Energies (IFAE), BIST, 08193 Bellaterra (Barcelona), Spain and Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig Lluís Companys 23, 08010 Barcelona, Spain”); D. Blixt (Scuola Superiore Meridionale, Italy) D.E. Bruschi (Institute for Quantum Computing Analytics (PGI-12), Forschungszentrum Jülich); L. Burderi (Department of Physics, University of Cagliari, Italy); L.I. Caramete (Institute of Space Science, Magurele, Romania); M. Chaichian (University of Helsinki, Finland); M. Chernyakova (Dublin City University); S. Clesse (Service de Physique Théorique, Université Libre de Bruxelles, Belgium); M. Demirci (Department of Physics, Karadeniz Technical University, Turkey); P.K. Dhankar (Department of Mathematics, G H Raisoni College of Engineering, Nagpur, India); F. Di Lodovico (King’s College London); T. Di Salvo (Department of Physics and Chemistry, University of Palermo, Italy); D.D. Dimitrijević (Faculty of Sciences and Mathematics, University of Niš, Serbia); J. Ellis (Physics Department, King’s College London, UK and Theoretical Physics Department, CERN, Switzerland); C. Escamilla-Rivera (Instituto de Ciencias Nucleares, Universidad Nacional Autónoma de México); L. Freidel (Perimeter Institute, Canada); A. Fuster (Department of Mathematics and Computer Science, Eindhoven University of Technology, The Netherlands); J. Gamboa (Universidad de Santiago de Chile); A. k. Ganguly (Banaras Hindu University, India); D.A. Gomes (University of Tartu, Estonia); D. Guetta (Physics Department, Ariel University, Israel); M.J. Guzman (University of Tartu, Estonia); F.J. Herranz (Departamento de Física, Universidad de Burgos, Spain); A. Ioannisian (ITPM and YerPhI, Yerevan 0036, Armenia); P. Jetzer (University of Zurich); K.H. Kampert (Department of Physics, University of Wuppertal, Germany); A. Karam (NICPB, Tallinn); F. Klinkhamer (Karlsruhe Institute of Technology); G. Lambiase (Università di Salerno, Italy); C. Lämmerzahl (ZARM, University of Bremen); J. Levi Said (University of Malta); L. Marchetti (University of New Brunswick, Canada); P. Martín-Moruno (Departamento de Física Teórica and IPARCOS, Universidad Complutense de Madrid, Spain); H. Martínez-Huerta (Departamento de Física y Matemáticas, Universidad de Monterrey, NL, México); M. Milošević (Faculty of Sciences and Mathematics, University of Niš, Serbia); L. Miramonti (University of Milano and INFN); J. Navarro-Salas (University of Valencia and IFIC, Spain); S. Navas (Dpto. de Fisica Teorica y del Cosmos and C.A.F.P.E., University of Granada, Spain); M. Niechciol (Experimentelle Astroteilchenphysik, Center for Particle Physics Siegen, University of Siegen, Germany); E.R. Nissimov (Insitute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Sofia, Bulgaria); R.C. Nunes (Instituto de Física, Universidade Federal do Rio Grande do Sul, Brazil); N.A. Obers (Niels Bohr Instistute and Nordita); G.J. Olmo (University of Valencia and IFIC, Spain); A. Ottewill (University College Dublin); A. Övgün (Physics Department, Eastern Mediterranean University, Turkey); C.F. Paganini (University of Regensburg, Germany); U. Pensec (Sorbonne Université, CNRS, France); C. Pérez de los Heros (Uppsala University); B. Puliçe (Sabancı University, İstanbul, Turkey); A. Racioppi (NICPB, Tallinn, Estonia); M.D. Rodríguez Frías (University of Alcala, Madrid); H. Sahlmann (Friedrich-Alexander-Universität Erlangen-Nürnberg, FAU); İ. Sakallı (Physics Department, Eastern Mediterranean University, Turkey); M. Sakellariadou (King’s College London); F. Schussler (IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France); G.H.W. Sigl (Universität Hamburg); J. Solà Peracaula (Universitat de Barcelona, UB); C.F. Sopuerta (Institute of Space Sciences, ICE, CSIC and IEEC); N. Stergioulas (Aristotle University of Thessaloniki); T.S. Stuttard (Niels Bohr Institute, Denmark); T. Thiemann (FAU Erlangen Nuremberg, Erlangen, Germany); T. Trześniewski (University of Wroclaw); C. Tzerefos (National and Kapodistrian University of Athens, Greece); D. Vernieri (University of Naples and INFN); N. Voicu (Transilvania University of Brasov, Romania); A. Wojnar (Department of Theoretical Physics and IPARCOS, Complutense University of Madrid, E-28040, Madrid, Spain); D. Zavrtanik (University of Nova Gorica, Slovenia) and J.D. Zornoza (IFIC, Univ. of Valencia - CSIC).

### v3

@@ Abstract
The unification of quantum mechanics and general relativity has long been elusive. Only recently have empirical predictions of various possible theories of quantum gravity been put to test, where a clear signal of quantum properties of gravity is still missing. The dawn of multi-messenger high-energy astrophysics has been tremendously beneficial, as it allows us to study particles with much higher energies and travelling much longer distances than possible in terrestrial experiments, but more progress is needed on several fronts.
@@ 1 Introduction
To bridge this gap between fundamental theories of QG and observations, a bottom-up approach emerged: QG phenomenology [9]. It is a framework that employs phenomenological models inspired by the results from the different fundamental approaches and by physical reasoning about QG, to identify effects which can be searched for in experiments. How and which of these effects emerge from the theory, depends on the theoretical model under consideration, for example: semi-classical field equations; the coupling between local tabletop low-energy quantum systems and gravity; modified gravity with additional scalar or vector fields or modified gravity with additional geometric fields such as torsion, non-metricity or non-linear connections; non-minimal couplings between gravity and matter; deviations from Lorentz invariance as in the standard model extension (SME) or in doubly special relativity (DSR); and many more [10, 11, 12, 13, 9, 14, 15, 16, 17]. These models help to devise the experimental setups required for searching for specific signatures, and the types of effects being sought. At the same time, evidence or bounds on any given effect are serving as guides for the construction of fundamental theories compatible with these phenomenological results.
@@
In this white paper the QG phenomenology community summarizes and highlights paths to advance and improve the search for QG effects. The overall purpose of this white paper is to outline the future phenomenological and experimental research of QG with astrophysical data.
We address some of the most promising observational windows to scales where we expect QG effects to set in, where we focus on selected effects caused by violations, deviations or modifications from Lorentz invariance, since these leave observables traces well below the Planck scale and are, if present, in reach of being either detected or falsified by multi-messenger astronomy observations in the near future. Moreover, we identify the necessary conditions for realistically boosting these searches with present and future facilities. Even more importantly, we aspire to stimulate developments of novel methods for searches for signatures of QG. Given that at this point there are no facilities designed for this specific purpose, and experimental tests are usually performed as side projects, we particularly hope to motivate a development of dedicated experiments, and that this white paper can be used as a road map in such endeavours. Surely, this white paper cannot cover all the aspects of the search for QG so we focus on selected promising prospects of QG searches in multi-messenger astronomy.
@@ 2.1.2 Some specific effects
A competing theoretical mechanism that can cause a cosmic vacuum refractive index is associated with certain exotic models of dark matter [51, 52]. They can be distinguished from spacetime foam models through the specific dependence of the refractive index on the photon frequency and the emergence or non-emergence of birefringence [53].
Birefringence (different propagation speeds for each light polarization) from QG spacetime with a polymer-like structure at microscales in loop quantum gravity [54]. Such effect can be caused by mass-dimension five operators in the effective field theory extension of the QED sector of the Standard Model [55], by a non-minimal coupling between the electromagnetic potential and curvature from renormalization of QED on curved spacetime [56], by the SME extension of the electromagnetic sector [57] or by premetric electrodynamics in general [58]. Birefringence effects are further discussed in section 3.2.1.
@@ 2.2 Time delays in the propagation of cosmic messengers
Time delays emerge already from the classical interaction between gravity and particles and fields propagating in vacuum that is described by the general relativistic geometry of spacetime. In this section, we briefly recall this classical achromatic vacuum time delay effects, that are already predicted by GR, before examining additional energy-dependent (generally four-momentum dependent) time delays, due to microscopic quantum fluctuations or foaminess of spacetime (see section 2.1), which are widely discussed in the literature of QG and its phenomenology. Then, we highlight the main challenges and difficulties in the search for such QG-induced time delays.
@@ 2.2.1 Classical time delays
Even in the absence of QG effects, the gravitational interaction causes vacuum time delays in the propagation of photons of any frequency and other cosmic messengers through spacetime, compared to their propagation in the absence of gravity. The most famous instance of this fact, which has been experimentally observed and actually constitutes one of the Solar-System tests of Einstein’s GR, is the Shapiro delay [89]: the observation that the travel time of a radar signal between two points in the vicinity of a gravitating object is increased compared to the travel time of the signal in the absence of the object. It is important to stress that the Shapiro delay is “achromatic”, in the sense that it does not depend on the energy or polarization of the photon. In general, we gather large amounts of information about the gravitational interaction based on the observation of cosmic messengers propagating through spacetime and eventually being detect by us. Thus, in the pertinent data analyses of telescopes observing cosmic messengers, it is of utmost importance to take into account the classical – and possibly the yet unknown additional QG-induced – vacuum time delay effects, in order to interpret the observations correctly. Famous examples where the classical time delays are already implemented in the data analysis are pulsar timing observations [90, 91] as well as cosmological redshift and distance observations [92]. Pictorially, the time delay effects can be understood in the sense that gravity acts as a certain effective refractive medium for light propagation, thus introducing an effective refractive index [93]. On the basis of Maxwell’s electrodynamics and Einstein’s GR, the predictions for the classical time delays are independent of the energy (frequency) and of the polarisation of the photons, as a consequence of the Einstein equivalence principle [94], which states that locally, that is, within an appropriately small region in the neighbourhood of each spacetime point, the weak equivalence principle and local Lorentz invariance holds, meaning that physics is described by special relativity (SR). In what follows we refer to this feature as local Lorentz symmetry. Due to the universal coupling with gravity, such GR-induced time delays do not only affect the propagation of light, but also all other standard model particles and fields including neutrinos [95] and, due to its self interaction, gravitational waves [96, 97].
The above examples from classical gravity clearly demonstrate that (achromatic and polarization independent) time delays are an important prediction of GR, which, on the one hand, allows us to test the theory, while on the other, gives us important information about the nature of the (classical) gravitational interaction, which is associated with the geometry of spacetime.
@@ 2.2.2 Possible quantum gravity-induced time delays
In addition to the general relativistic, achromatic and polarisation independent time delays discussed so far, time delays resulting from purely quantum effects of gravity might emerge. In contrast to the classical effects, they might depend on the energy and the polarisation of the probe, for the the following reason. Highly energetic particles and fields will interact with gravity on smaller length scales than low energetic ones. It is expected that the smaller the length scale which the particles and fields probe, i.e. the higher their energy, the more the particles and fields are affected by quantum vacuum fluctuations of gravity or spacetime foam, see also the discussion in section 2.1. Effectively, this scenario can be described as particles and fields propagating through a (refractive) medium generated by vacuum fluctuations of gravity, capturing aspects of a quantum spacetime or spacetime foam. The mathematical description of this pictorial idea has been further developed within various formal and phenomenological frameworks of QG.
@@ 2.2.3 Theoretical challenges
A particularly important challenge is the distinction between violation and deformation of local Lorentz invariance. In other words, either the LIV implies the existence of a preferred reference frame, or the Lorentz group is replaced by the notion of spacetime and observer symmetry (DSR). Both, the absence or the detection of any kind of time delay can be used as a guiding principle to identify the scale and the type of deviations from local Lorentz symmetry. Whether this happens effectively at an intermediate length/energy scale or at the fundamental QG scale is unclear and depends on the theoretical approach under consideration. We would like to point out that for effects below the natural QG scale it is also important to distinguish between QG time delays and time delays due to a possible non-minimal coupling between curvature and electromagnetism, that might emerge due to renormalisation of QED on curved spacetime [56].
@@
In summary, it is known that gravity causes certain types of vacuum time delays. The search for additional, yet undetected energy-dependent time delays caused by potential QG effects can be an important step in finding evidences of a theory of QG. Thus, it is crucial to devise future dedicated experiments to address such a possibility both in an astrophysical context and in that of the more controllable Earth-based experiments. These two regimes complement each other, in the sense that they probe QG time delays on different scales, which have different systematic uncertainties.
@@ 2.3 Modifications of the kinematics of interactions of cosmic messengers
Threshold anomalies can be seen as kinematic modifications of the SR description of interactions, and indeed, it is usually the only effect that is modelled to estimate the phenomenological consequences of LIV or DSR on the interactions of cosmic messengers [114, 101]. However, the decay rates or cross sections of a process in LIV or DSR scenarios may differ from the standard ones even at energy ranges where the process is allowed in both, SR and its modifications, which may be important for effects such as the possible modification of the transparency of the universe to high-energy gamma rays. The LIV scenario has been explored in the past, but the availability of results for cross sections or decay rates is limited [115, 116, 117]. For the DSR scenario, changes in the cross sections or decay rates may only play a role in the case of a QG energy scale much lower than the Planck one, comparable to the energies of cosmic messengers. While such dynamical calculations can be carried our for the LIV case in the framework of effective field theory, the DSR scenario still lacks a well-defined dynamical approach. In fact, an appropriate framework for DSR should also address some of its known theoretical ambiguities, such as the soccer-ball problem [118, 119, 120, 121, 122], the spectator problem [123, 124, 125, 126], or issues with momentum labelling [127, 122, 128], whose resolution may require a novel perspective on DSR [129]. Consequently, the implications of threshold anomalies in DSR remain largely unexplored, for which it will be necessary to develop new approaches and techniques [113].
@@ 2.3.1 Hadronic processes (affecting cosmic rays)
UHECR at the highest energies can undergo interactions with background photons, such as pion photoproduction and (for nuclei heavier than protons) photodisintegration. This introduces interaction horizons, i.e. a maximum distance from which cosmic rays above a given energy can be observed. Moreover, such interactions may produce fluxes of cosmogenic neutrinos and photons. Both these effects can be altered by LIV in the hadronic sector [130, 131, 132, 133, 134], so that UHECR observations could in principle be used to test LIV scenarios. Unfortunately, as discussed in section 3.1, several uncertainties about the sources and propagation of UHECRs complicate the interpretation of the UHECR energy spectrum and mass composition at Earth in terms of those escaping their source environment. Furthermore, the most recent such studies [135, 136, 137, 138] suggest that the spectral shape at the sources is harder than what is predicted by the Fermi mechanism, and that at the highest energies UHECRs are increasingly heavy nuclei, escaping their sources with Lorentz factors near or below the threshold for pion photoproduction. This substantially reduces their sensitivity to LIV effects, as well as the production rates of cosmogenic neutrino and photons, compared to the case of protons with the same energy. On the other hand, the presence of a non-negligible proton fraction at the highest energies can still not be fully ruled out by the most recent data, and it could be tested in the future with the aim of increasing the constraining power of the UHECR spectrum and composition data, as will be mentioned in section 4.3.
@@ 3.2.1 Propagation effects in vacuum
The temporal distribution of emitted particles is usually described with a light curve template defined by several parameters. The magnitude of uncertainties of these parameters varies from case to case, and is a matter of compromise between several factors. Faster and more pronounced flux variability puts stronger constraints on the emission time: the sensitivity to the QG energy scale is inversely proportional to the variability time scale for linear () QG effects and to its square root for quadratic () ones [101, Table 1]. This, however, requires using shorter time bins, which results in relatively smaller number of events per bin, and, consequently, larger uncertainties. For this reason, brighter, more variable emissions are selected to study time delays. Improvement of instrumental sensitivity will directly reflect on the ability to model the emission light curves, and on the sensitivity to the QG energy scale.
@@
LIV effects could introduce modifications to the phase and group velocities of the circularly polarized modes of photons. When these modes have different velocities, propagation of light is subject to vacuum birefringence, which induces a rotation of the polarization direction of linearly polarized signals, or a loss of polarization degree [224, 225, 226, 227]. To observe this effect, knowledge of the emission properties of the photon source is required. Moreover, since the degree of polarization directly depends on propagation details, the properties of the medium ought to be known. For instance, uncertainties in the magnetic field could influence the polarization through Faraday rotation.
@@ 4.3 Cosmic-ray experiments
Presently our most powerful cosmic-ray telescope is the Pierre Auger Observatory, and its results have provided the basis for several QG-phenomenology analyses [259, and references therein]. However, in general converting UHECR data into constraints on QG parameters faces several challenges due to our lack of knowledge about the sources and propagation of UHECR, described in section 3.1, but one can be optimistic about progress in these directions in the coming years [198].
@@ 6 Final remarks
The scale of is therefore one of the most pressing theoretical unknowns, and experimental efforts in QG phenomenology should aim for the identification of this scale. Unfortunately, given the current understanding of the consequences of a QG theory, it is not possible to offer very specific proposals to achieve this objective. As a generic goal, this corresponds to trying to push the high-energy limit in the present and near-future experiments. This may be crucial for the observation of sharp cutoffs in the spectrum of photons or neutrinos predicted in superluminal LIV models, as well as to observe an increase of the transparency of the universe for high-energy photons or cosmic rays in the case of DSR or subluminal LIV models. Another relevant effect that may be observed with better exposition at high energies is an enhancement in the cosmogenic production of neutrinos and photons by the interactions of cosmic rays with the CMB and EBL.
@@
A.F. Ali (Essex County College, NJ, USA); M. Arzano (University of Naples and INFN); J-L. Atteia (IRAP, Université de Toulouse, CNRS, CNES, UPS, Toulouse, France); C. Barceló (Instituto de Astrofísica de Andalucía, IAA CSIC); I. Ben-Dayan; D. Benisty (University of Cambridge, and Frankfurt Institute for Advanced Studies, fias); N. Bilic (Rudjer Boskovic Institute, Zagreb, Croatia); O. Birnholtz (Bar Ilan University, Israel); D. Blas (”Institut de Física d’Altes Energies (IFAE), BIST, 08193 Bellaterra (Barcelona), Spain and Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig Lluís Companys 23, 08010 Barcelona, Spain”); D. Blixt (Scuola Superiore Meridionale, Italy) D.E. Bruschi (Institute for Quantum Computing Analytics (PGI-12), Forschungszentrum Jülich); L. Burderi (Department of Physics, University of Cagliari, Italy); L.I. Caramete (Institute of Space Science, Magurele, Romania); M. Chaichian (University of Helsinki, Finland); M. Chernyakova (Dublin City University); S. Clesse (Service de Physique Théorique, Université Libre de Bruxelles, Belgium); M. Demirci (Department of Physics, Karadeniz Technical University, Turkey); P.K. Dhankar (Department of Mathematics, G H Raisoni College of Engineering, Nagpur, India); F. Di Lodovico (King’s College London); T. Di Salvo (Department of Physics and Chemistry, University of Palermo, Italy); D.D. Dimitrijević (Faculty of Sciences and Mathematics, University of Niš, Serbia); J. Ellis (Physics Department, King’s College London, UK and Theoretical Physics Department, CERN, Switzerland); C. Escamilla-Rivera (Instituto de Ciencias Nucleares, Universidad Nacional Autónoma de México); L. Freidel (Perimeter Institute, Canada); A. Fuster (Department of Mathematics and Computer Science, Eindhoven University of Technology, The Netherlands); J. Gamboa (Universidad de Santiago de Chile); A. k. Ganguly (Banaras Hindu University, India); D.A. Gomes (University of Tartu, Estonia); M.J. Guzman (University of Tartu, Estonia); F.J. Herranz (Departamento de Física, Universidad de Burgos, Spain); A. Ioannisian (ITPM and YerPhI, Yerevan 0036, Armenia); P. Jetzer (University of Zurich); K.H. Kampert (Department of Physics, University of Wuppertal, Germany); A. Karam (NICPB, Tallinn); F. Klinkhamer (Karlsruhe Institute of Technology); G. Lambiase (Università di Salerno, Italy); C. Lämmerzahl (ZARM, University of Bremen, Germany); J. Levi Said (University of Malta); M. D. Maia, (Institute of Physics, University of Brasília, Brasília D.F., Brazil); L. Marchetti (University of New Brunswick, Canada); P. Martín-Moruno (Departamento de Física Teórica and IPARCOS, Universidad Complutense de Madrid, Spain); H. Martínez-Huerta (Departamento de Física y Matemáticas, Universidad de Monterrey, NL, México); M. Milošević (Faculty of Sciences and Mathematics, University of Niš, Serbia); L. Miramonti (University of Milano and INFN); J. Navarro-Salas (University of Valencia and IFIC, Spain); S. Navas (Dpto. de Fisica Teorica y del Cosmos and C.A.F.P.E., University of Granada, Spain); M. Niechciol (Experimentelle Astroteilchenphysik, Center for Particle Physics Siegen, University of Siegen, Germany); E.R. Nissimov (Insitute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Sofia, Bulgaria); R.C. Nunes (Instituto de Física, Universidade Federal do Rio Grande do Sul, Brazil); N.A. Obers (Niels Bohr Instistute and Nordita); G.J. Olmo (University of Valencia and IFIC, Spain); A. Ottewill (University College Dublin); A. Övgün (Physics Department, Eastern Mediterranean University, Turkey); C.F. Paganini (University of Regensburg, Germany); U. Pensec (Sorbonne Université, CNRS, France); C. Pérez de los Heros (Uppsala University); B. Puliçe (Sabancı University, İstanbul, Turkey); A. Racioppi (NICPB, Tallinn, Estonia); M.D. Rodríguez Frías (University of Alcala, Madrid); H. Sahlmann (Friedrich-Alexander-Universität Erlangen-Nürnberg, FAU); İ. Sakallı (Physics Department, Eastern Mediterranean University, Turkey); M. Sakellariadou (King’s College London); F. Schussler (IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France); G.H.W. Sigl (Universität Hamburg); J. Solà Peracaula (Universitat de Barcelona, UB); C.F. Sopuerta (Institute of Space Sciences, ICE, CSIC and IEEC); N. Stergioulas (Aristotle University of Thessaloniki); T.S. Stuttard (Niels Bohr Institute, Denmark); T. Thiemann (FAU Erlangen Nuremberg, Erlangen, Germany); T. Trześniewski (University of Wroclaw); C. Tzerefos (National and Kapodistrian University of Athens, Greece); D. Vernieri (University of Naples and INFN); N. Voicu (Transilvania University of Brasov, Romania); A. Wojnar (Department of Theoretical Physics and IPARCOS, Complutense University of Madrid, E-28040, Madrid, Spain); D. Zavrtanik (University of Nova Gorica, Slovenia) and J.D. Zornoza (IFIC, Univ. of Valencia - CSIC).